OpenStudy (anonymous):

prove cos4A = 8sin^4(a) - 8sin^2(A) + 1 i've tried all sorts of methods but i just can't get the answer

OpenStudy (anonymous):

@ganeshie8

OpenStudy (gorv):

is it cos (4a) ???

OpenStudy (gorv):

cos2x=2cos^2x-1

OpenStudy (gorv):

cos(2*(2a)) x=2a here

OpenStudy (gorv):

cos (2*(2a))=2cos^2(2a)-1

OpenStudy (gorv):

\[2(\cos2a)^2-1\]

OpenStudy (gorv):

cos2a=2cos ^2a-1

OpenStudy (gorv):

=2(2cos^2a-1)^2-1

OpenStudy (anonymous):

yeah it is hang on i'll be looking through your working

OpenStudy (anonymous):

but then it becomes 2(4cos^4A-1)^2-1 which i can't expand because it'll be expanding a power 4..

OpenStudy (gorv):

you dont how to open the square

OpenStudy (gorv):

cos^a=1-sin^a

OpenStudy (gorv):

2(2(1-sin^2a)-1)^2-1

OpenStudy (gorv):

=2(2-2sin^2a-1)^2-1 =2(1-2sin^2a)^2-1

OpenStudy (gorv):

2(1^2+(2sin^2a)^2-2*1*2sin^2a)-1 =2(1+4sin^4(a)-4sin^2a)-1 =2+8sin^4a-8sin^2a-1 =8sin^4a-8sin^2a+1

OpenStudy (gorv):

one thing i would like to say frn that if someone help u out thn its your responsibility to rply yes/no/ or anythimg else..not to sit like u r watching tv

OpenStudy (anonymous):

well, thing is, you're posting every step one by one and its kind of infuriating because you can actually do them in fewer posts i'm trying to process with information with each step, that's why i don't reply. and since you're posting in that sequence for some reason —nobody else does that— i decided to just wait till you were done but thank you for helping